# Tuning Large Multimodal Models for Videos using Reinforcement Learning from AI Feedback

### v1

@@ Abstract
Recent advancements in large language models have influenced the development of video large multimodal models (VLMMs). The previous approaches for VLMMs involved Supervised Fine-Tuning (SFT) with instruction-tuned datasets, integrating LLM with visual encoders, and adding additional learnable modules. Video and text multimodal alignment remains challenging, primarily due to the deficient volume and quality of multimodal instruction-tune data compared to text-only data. We present a novel alignment strategy that employs multimodal AI system to oversee itself called Reinforcement Learning from AI Feedback (RLAIF), providing self-preference feedback to refine itself and facilitating the alignment of video and text modalities. In specific, we propose context-aware reward modeling by providing detailed video descriptions as context during the generation of preference feedback in order to enrich the understanding of video content. Demonstrating enhanced performance across diverse video benchmarks, our multimodal RLAIF approach, VLM-RLAIF, outperforms existing approaches, including the SFT model. We commit to open-sourcing our code, models, and datasets to foster further research in this area. https://github.com/yonseivnl/vlm-rlaif
Figure 1: Qualitative comparison of VLMMs on various video benchmarks. The video question answering task is marked in purple, video-based generative task in blue, the text-to-video retrieval task in violet and the action recognition task in orange color. VLM-RLAIF achieves superior performances on a broad range of video benchmarks compared to previous approaches, including VLM-SFT.
Figure 2: Illustration of the proposed VLM-RLAIF. LLM tuned with video-text instruction-following data, i.e., multimodal SFT model, often produces responses that are not temporally and visually grounded to the video input, as depicted in red color. We propose a method that involves using the VLMM to supervise itself by providing self-preference feedback of generated responses, refining itself and facilitating the alignment of video and text modalities.
@@ 1 Introduction
Recent progress in large language models (LLMs) has advanced the field of video large multimodal models (VLMMs) Li et al. (2023a); Muhammad Maaz and Khan (2023); Lin et al. (2023). By extending the logical reasoning and advanced cognitive capabilities of LLMs to the visual domain, these advancements have enabled VLMMs to showcase remarkable proficiency in tasks such as video understanding Li et al. (2023a), video question answering Ko et al. (2023) and instruction-following tasks Muhammad Maaz and Khan (2023); Luo et al. (2023).
@@
These VLMMs achieve this by integrating pre-trained LLMs, such as LLaMA Touvron et al. (2023), with visual encoders and an additional learnable module, i.e., LoRA Hu et al. (2022). Specifically, they undergo a supervised fine-tuning (SFT) stage using multimodal instruction-tune data Luo et al. (2023); Muhammad Maaz and Khan (2023); Li et al. (2023a). This process aims to adapt LLMs to video modality, thus improving their ability to interpret visual content.
@@
However, there is a challenge in achieving seamless multimodal alignment between video and text, primarily due to the deficiency in volume and quality of multimodal instruction-tune data compared to text-only data. Text-only data are typically abundant and diverse, while multimodal data are often limited in both quantity and comprehensiveness Wei et al. (2021); Liu et al. (2023a). This often leads the VLMMs to generate responses that are not properly grounded in the visual context, as illustrated by the SFT model responses in Fig. 2.
@@
To address this, we propose a novel alignment approach for video and text modality that involves using the VLMM to supervise itself by providing preference feedback of generated responses, as shown in Fig. 2. Specifically, we introduce Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022); Lee et al. (2023) for multimodal alignment. Unlike Reinforcement Learning with Human Feedback (RLHF), which has been successful in aligning text-only or image-text based AI agents with human intentions Ouyang et al. (2022); Sun et al. (2023a); Yu et al. (2023), the RLAIF allows for scalable oversight with minimum human intervention. Particularly, we enhance AI’s feedback by proposing a context-aware reward modeling, focusing on improved clarity in video content (Sec. 3.2). Furthermore, to compensate for the limited multimodal instruction-tune data for training the SFT, we augment it with a human-labeled video question answering and an object-centric multimodal instruction-tune dataset. Further, to effectively utilize this expanded dataset, we propose a simple curriculum training strategy for enhancing the alignment between video and text modality (Sec. 3.3).
@@
To sum up, we call our proposed framework of training an VLMM with AI feedback as Video Large Multimodal model with RLAIF or VLM-RLAIF for short. Our empirical studies show that our aligned VLM-RLAIF exhibits superior performance compared to state-of-the-art VLMMs across a wide array of video benchmarks, as illustrated in Fig. 1.
@@ 2 Related Work
In recent research efforts, there has been a significant focus on advancing Language Models (LLMs) to go beyond natural language understanding, extending into the realm of multi-modal comprehension. The goal is to develop LLMs capable of interpret additional modalities, e.g., image Liu et al. (2023a), video Li et al. (2023a); Lin et al. (2023), 3D point-cloud Guo et al. (2023), etc.
To provide additional modality into LLMs, most of work utilize pre-trained encoder, such as CLIP Radford et al. (2021), Q-former Li et al. (2022) or ImageBind Girdhar et al. (2023), to extract each modality feature from data. These features then projected to the token embedding space of the language model. Following this, the models are subject to supervised fine-tuning (SFT) using synthetically generated, modality-specific instruction-following datasets. These approaches, exemplified in works such as LLaVA Liu et al. (2023a), Video-LLaVA Lin et al. (2023) or Point-LLM Guo et al. (2023), facilitates the development of models proficient in engaging in conversations grounded in additional modality.
Figure 3: Detailed illustration of the three stages for context-aware reward modeling. Dotted box in each stage represents distinct model input: the first is for generating responses using the SFT model, the second for the judge model to evaluate and choose the superior response between options A and B, and the third for training the RM. Each input includes a task-specific prompt, denoted by the yellow box, tailored to guide the model’s operation within its respective function (more details are in Appendix Sec. C). The operational flow of each stage is as follows: (1) The SFT model produces two candidate responses from the provided video and question. (2) With the video, question and responses at hand, the SFT model utilize context information and guiding prompt to evaluate the responses. (3) The RM is trained using the preference pairs generated in the previous step as indicated in orange box.
Figure 4: An example of context-aware preference selection. We demonstrate our model’s ability to generate preference feedback, i.e., preferred choice and explanation marked in orange box, on given instruction input, two responses and context. Red color denotes an incorrect response, while blue color signifies a correctly grounded response with respect to the visual input. The rationale behind selecting ‘Output B’ as the preferred choice is indicated in green.
@@
To operate the model safely and in accordance with human intentions, Reinforcement Learning from Human Feedback (RLHF) has been proposed as a viable solution Ouyang et al. (2022); Sun et al. (2023a). By collecting the preferences from human evaluators, it usually trains the reward model that gives a high reward to the preferred output of the model. However, significant challenge in this process is the annotation cost associated with selecting preference. To mitigate this issue, the concept of Reinforcement Learning with AI-inferred Feedback (RLAIF) was introduced Bai et al. (2022); Lee et al. (2023); Sun et al. (2023b). RLAIF capitalizes on the inherent ability of Large Language Models (LLMs) to rate or evaluate information, allowing the LLM itself to assign preferences.
@@ 3 VLM-RLAIF Framework
3.1 Overview
Contrasting to the limited scalability of human feedback in RLHF, we leverage AI’s feedback for aligning multimodality between video and text, reducing the reliance on exhaustive human-annotated preferences Ouyang et al. (2022); Sun et al. (2023a). The training procedure of VLM-RLAIF can be summarized into three stages as follows:
@@ 3.1 Overview
We fine-tune an LLM, i.e., Vicuna, using supervised learning on synthetically generated video-text instruction-tune data Muhammad Maaz and Khan (2023). This involves integrating a vision encoder with two linear layers and an additional learnable parameters using LoRA Hu et al. (2022), into the training process. Such fine-tuning enhances the model’s capability to follow instructions. We refer to this model as the Video Large Multimodal model with SFT or VLM-SFT for short.
@@
A key aspect of the RLAIF involves leveraging a pre-trained AI model to generate human-like preferences between different responses generated from the same input Bai et al. (2022); Sun et al. (2023b); Lee et al. (2023). To accomplish this, we employ the VLM-SFT as a judge for assessing preferences. Once preferences are annotated, we proceed to train a reward model (RM) based on these annotated labels using a cross-entropy loss, following the Bradley-Terry model for estimating score functions from pairwise preferences Ouyang et al. (2022); Sun et al. (2023a). We illustrate the training procedure for collecting preferences and training the reward model in the Fig. 3. The RM give higher score reward to the better response and lower score reward to the less appropriate one in a pair of responses, thus providing guidance to a policy model during the reinforcement learning (RL) process.
Reinforcement learning.
During the RL phase, we fine-tune a supervised policy model, initialized from the VLM-SFT, aiming to optimize the scalar reward output from the trained RM. We use the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) for this purpose, following prior research Ouyang et al. (2022); Sun et al. (2023b, a). This process finely tunes the model to respond optimally to the rewards, thereby enhancing its decision-making abilities in accordance with the learned preferences.
@@ 3.2 Context-Aware Reward Modeling
For effective utilization of VLM-SFT in preference selection, a nuanced understanding of video content is essential, particularly in scenarios involving complex scenes with elements subject to varied interpretations. Acknowledging this intricacy, it becomes apparent that exclusive dependence on VLM-SFT for direct video analysis may fall short in achieving thorough comprehension.
@@
In response, we enhance the VLM-SFT by integrating detailed video descriptions, referred to as context, into the preference selection workflow, thereby imparting additional contextual clarity, as illustrated in Figures 3-(2) and 4. This process entails breaking down the video into smaller clips, followed by generating descriptive caption for each segment to obtain fine-grained information. Subsequently, these individual captions are concatenated into an exhaustive narrative of the video. This detailed narrative is then provided to the judge model, i.e., VLM-SFT, to aid in better preference selection. This process not only improves the VLM-SFT’s ability to obtain a comprehensive view of the video content, but also enables it to more identify the most suitable response. Integrating all inputs, i.e., marked in dotted box in Fig. 3-(2), with a specific prompt, i.e., rules for generating preferences as illustrated in Appendix Fig. 9, allows us to collect context-aware synthetic preferences, thus enriching the model’s decision-making process.
Training reward model.
The reward model (RM) is designed to assign higher scores to responses considered better and lower scores to those considered lower in quality. Starting from VLM-SFT with the final linear layer removed, we train a model to take in an input prompt and response, i.e., marked in dotted box of Fig. 3-(3), and output a scalar reward value. Using the preference dataset produced from the VLM-SFT, we train the RM with a cross-entropy loss. Specifically, we use 13B VLM-SFT to train the reward model, as it gives slightly better performance than 7B VLM-SFT (see the quantitative comparison discussed in Appendix Sec. A). Remarkably, after training RLAIF using the RM, the 7B VLM-RLAIF significantly surpass the 13B VLM-SFT, thus validating the effectiveness of our proposed framework in aligning video and text modalities (refer to Sec. 4.2 for more details).
3.3 Two-stage Curriculum Supervised Fine-Tuning
To obtain video conversational capability, we train the LLM with a open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023). Further, to enhance this, we not only expand our training with additional video-text instruction-tune datasets but also introduce a simple curriculum learning strategy. This approach is designed to improve the model’s performance progressively, ensuring a more effective and nuanced understanding of video-text interactions.
@@ 3.3 Two-stage Curriculum Supervised Fine-Tuning
To improve the video understanding ability, we first augment the open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023) with existing human-annotated video question answering datasets Xiao et al. (2021); Li et al. (2020). Moreover, recognizing the importance of aligning LLMs with the video modality, we focus on obtaining instruction-tune dataset that encompass both visual and temporal intricacies. To achieve this, we leverage Video Localized Narratives Voigtlaender et al. (2023), a dataset that provides detailed descriptions of the appearance and complex actions of all objects in the video. We transform the narrative dataset into object-centric instruction-tune dataset format using ChatGPT similar to prior work Muhammad Maaz and Khan (2023). More details about the generated instruction-tune dataset can be found in the Appendix Sec. B. We train the SFT model using both the open-sourced and our proposed datasets, resulting in improved performance, as discussed in the Sec. 4.3.
Curriculum supervised fine-tuning
Recent VLMM research has advanced instruction-following capabilities through curating multimodal instruction-tune data Muhammad Maaz and Khan (2023); Luo et al. (2023), joint learning across image and video datasets Lin et al. (2023); Li et al. (2023c), and enhanced bridging networks like adapter Liu et al. (2023b). Unlike these approaches, we emphasize a structured learning method for VLMMs, inspired by the human learning process. We advocate for the importance of the training sequence, drawing from curriculum learning principles Chang et al. (2021), and propose a simple two-stage curriculum learning strategy for SFT, aiming to optimize the instruction-following proficiency of VLMMs.
To accomplish this, we divide the SFT dataset based on complexity into two segments: ‘easy’ and ‘hard’. The model is first trained on ‘easy’ tasks to learn basic concepts, then progressed to ‘hard’ tasks for advanced challenges. Given that longer answers often require deeper comprehension of the context and enhanced proficiency in creating syntactically complex sentences Xu et al. (2020); Agrawal and Singh (2023); Ranaldi et al. (2023), we use answer length as our proxy for sample difficulty, meaning that the longer the answer sentence, the more difficult the task is. This proxy also allows for a clear basis for dividing data, offering a concrete measure of difficulty that aligns the model’s learning trajectory with the incremental nature of human learning.
@@
Correctness Detail Context Temporal Consistency
VideoChat Li et al. (2023a) 7B 2.23 2.50 2.53 1.94 2.24
@@
Table 1: Qualitative comparison between different VLMMs on video-based generative performance benchmark. We assess various VLMMs on video-based generative benchmark which measures various generative aspects. Our approach, VLM-RLAIF, shows a performance improvement over previous approaches, with the exception of GPT-4V, and demonstrates noticeable enhancements across five criteria when compared to the VLM-SFT. Here, $\Delta$ (RLAIF - SFT) indicates the improvement of RLAIF model over SFT model. $\dagger$} denotes reproduced results using the author’s implementation.

### v3

@@ Abstract
Recent advancements in large language models have influenced the development of video large multimodal models (VLMMs). Previous approaches for VLMMs involve Supervised Fine-Tuning (SFT) with instruction-tuned datasets, integrating LLM with visual encoders, and additional learnable parameters. Here, aligning video with text, and vice versa, remains a challenge, primarily due to the insufficient quality and quantity of multimodal instruction-tune data compared to that of text-only. This discrepancy often results in alignments that poorly ground the video content. To address this, we present a novel alignment strategy that employs a multimodal AI system equipped with Reinforcement Learning from AI Feedback (RLAIF), providing self-preference feedback to refine itself and facilitating the alignment of video and text modalities. Our approach uniquely integrates detailed video descriptions as context into a multimodal AI system during preference feedback generation to enrich the understanding of video content, a process we call context-aware reward modeling. Empirical evaluations on various video benchmarks demonstrate that our VLM-RLAIF outperforms existing approaches, including the SFT model. We commit to open-sourcing our code, models, and datasets to foster further research in this area. https://github.com/yonseivnl/vlm-rlaif
Figure 1: Quantitative comparison of VLMMs on various video benchmarks. The video question answering (VQA) task is marked in purple, video-based generative task in blue, the text-to-video (T2V) retrieval task in violet and the action recognition (AR) task in orange color. VLM-RLAIF achieves superior performances on a broad range of video benchmarks compared to previous approaches, including VLM-SFT. Comprehensive comparisons are provided in Tables 1, 2 and 3.
@@ 1 Introduction
Large language models (LLMs) are advancing many language and multimodal AI tasks, including those involved with video large multimodal models (VLMMs) Li et al. (2023b); Muhammad Maaz and Khan (2023); Lin et al. (2023). Extending the logical reasoning and advanced cognitive capabilities of LLMs to the visual domain, VLMMs are now remarkably proficient in tasks such as video understanding Li et al. (2023b), video question answering Ko et al. (2023) and instruction-following tasks Muhammad Maaz and Khan (2023); Luo et al. (2023). These models include publicly available LLMs Touvron et al. (2023); Chiang et al. (2023); Taori et al. (2023) with visual encoders and additional learnable parameters Hu et al. (2022); Liu et al. (2023b); Li et al. (2023a). To adapt LLMs to the video modality, thus improving their ability to interpret visual content, they all undergo a supervised fine-tuning (SFT) stage using multimodal instruction-tune data Luo et al. (2023); Muhammad Maaz and Khan (2023); Li et al. (2023b).
Figure 2: Illustration of the proposed VLM-RLAIF. An LLM tuned with video-text instruction-following data, i.e., multimodal SFT model, often produces responses that are not temporally and visually grounded to the video input, as depicted in red color. We propose a method that involves using the VLMM to supervise itself by providing self-preference feedback of generated responses using reward model, refining itself and facilitating the alignment of video and text modalities.
@@
However, multimodal alignment between video and text faces a significant challenge of deficiency in volume and quality of multimodal instruction-tune data compared to text-only data; text-only data are typically abundant and diverse, while multimodal data are often limited in both quantity and comprehensiveness Wei et al. (2021); Liu et al. (2023a). This often leads the VLMMs to generate responses that are not properly grounded in the visual content, as illustrated by the SFT model responses in Fig. 2.
@@
To address the issue of VLMMs producing inadequately grounded response, we propose a novel method to align video with text that involves using the VLMM to supervise itself by providing preference feedback of generated responses, as shown in Fig. 2. Specifically, we propose to use Reinforcement Learning from AI Feedback (RLAIF) Bai et al. (2022); Lee et al. (2023) for multimodal alignment. Unlike the Reinforcement Learning with Human Feedback (RLHF), which has been successful in aligning text-only or image-text based AI agents with human intentions Ouyang et al. (2022); Sun et al. (2023a); Yu et al. (2023), the RLAIF allows for scalable oversight with minimal human intervention. In particular, we improve AI feedback by leveraging detailed video descriptions as a context during the generation of preference feedback, focusing on improved clarity in video content, a process we refer to as context-aware reward modeling (Sec. 3.1). In addition, to compensate for the limited multimodal instruction-tune data for training the SFT, we augment it with a human-labeled video question answering and an object-centric multimodal instruction-tune dataset. Further, to effectively utilize this expanded dataset, we propose a simple curriculum training strategy for enhancing the alignment between video and text modality (Sec. 3.2).
@@
We call our proposed framework of training an VLMM with AI feedback as Video Large Multimodal model with RLAIF or VLM-RLAIF for short. Our empirical studies show that our aligned VLM-RLAIF exhibits superior performance compared to state-of-the-art VLMMs across a wide array of video benchmarks, as illustrated in Fig. 1.
@@ 2 Related Work
Recently, there have been significant advances for LLMs to go beyond natural language understanding, extending into the realm of multimodal comprehension. The goal is to develop LLMs capable of understanding various modalities, e.g., image Liu et al. (2023a), video Li et al. (2023b); Lin et al. (2023), 3D point-cloud Guo et al. (2023) and etc.
To make the LLMs multimodal, most of the work utilize a pretrained encoder, such as CLIP Radford et al. (2021a), Q-former Li et al. (2022) or ImageBind Girdhar et al. (2023), to extract each modality’s representations from data. These representations are then projected into the token embedding space of the language model. Then, the models undergo supervised fine-tuning (SFT) with synthetically generated, modality-specific instruction-following datasets. These approaches, adopted in LLaVA Liu et al. (2023a), Video-LLaVA Lin et al. (2023) or Point-LLM Guo et al. (2023), facilitate the development of proficient conversations grounded in additional modality.
@@
To operate the model safely and in accordance with human intentions, Reinforcement Learning from Human Feedback (RLHF) has been proposed as a viable solution Ouyang et al. (2022); Sun et al. (2023a). By collecting preferences from human evaluators, it usually trains the reward model that gives a high reward to the preferred output of the model. However, a significant challenge in this process is the annotation cost associated with selecting the preference. To mitigate this issue, Reinforcement Learning from AI Feedback (RLAIF) was proposed Bai et al. (2022); Lee et al. (2023); Sun et al. (2023b). RLAIF capitalizes on the inherent ability of Large Language Models (LLMs) to evaluate the generated responses from the SFT model, allowing the LLM itself to assign preferences.
@@ 3 VLM-RLAIF Framework
To overcome the limited scalability of human feedback in RLHF, we use AI’s feedback to align multimodality between video and text, reducing the reliance on exhaustive human-annotated preferences Ouyang et al. (2022); Sun et al. (2023a). In particular, we improve the feedback process by using detailed video descriptions, thereby achieving better contextual clarity in video content. The training procedure of VLM-RLAIF can be summarized into three stages as follows:
@@
We first fine-tune an LLM, e.g., Vicuna, using supervised learning on synthetically generated video-text instruction-tune data Muhammad Maaz and Khan (2023). This involves the integration of a vision encoder with two linear layers and additional learnable parameters using LoRA Hu et al. (2022), into the training process. This fine-tuning allows the model to better follow the instructions Muhammad Maaz and Khan (2023); Su et al. (2023). Additionally, we improve the SFT process by expanding the instruction-tune data and introducing simple curriculum learning (Sec. 3.2). We refer to this fine-tuned model as the Video Large Multimodal model with SFT or VLM-SFT for short.
@@
A key aspect of the RLAIF involves leveraging a pre-trained AI model to generate human-like preferences between different responses generated from the same input Bai et al. (2022); Sun et al. (2023b); Lee et al. (2023). To obtain human-like preference, we employ the VLM-SFT as a judge to assess preferences. Once preferences are judged, we train a reward model (RM) based on preferences using a cross-entropy loss, following the Bradley-Terry model for estimating score functions from pairwise preferences Ouyang et al. (2022); Sun et al. (2023a). We describe the training procedure for collecting preferences and training the reward model in Sec. 3.1. The RM give higher score reward to the better response and lower score reward to the less appropriate one in a pair of responses (see examples in Appendix Fig. 12), thus guiding the policy model using reinforcement learning (RL).
Reinforcement learning from AI feedback.
We finally fine-tune a supervised policy model, initialized from the VLM-SFT, aiming to optimize the scalar reward output of the trained RM by reinforcement learning. Specifically, we use the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017), following Ouyang et al. (2022); Sun et al. (2023b, a).
@@ 3.1 Context-Aware Reward Modeling
For VLM-SFT to select preference grounded on the video, we argue that a detailed understanding of video content is necessary for more accurate and contextually relevant decisions by the VLM-SFT. However, the current video encoder presents challenges in accurately encoding the temporal details of videos as they are based on the image encoder Radford et al. (2021b).
Figure 3: The three stages of the proposed context-aware reward modeling. The work flow of each stage is as follows: (1) The SFT model produces two candidate responses from the provided video and question. (2) With the video, question and responses at hand, the SFT model utilize context information and guiding prompt to evaluate the responses. (3) The RM is trained using the preference pairs generated in the previous step as indicated in orange box. Dotted box in each stage denotes a model’s input: the first is for generating responses using the SFT model, the second is for the judge model to evaluate and choose the superior response between options A and B, and the third is for training the RM. Each input includes a task-specific prompt, denoted by the yellow box, tailored to guide the model’s operation within its respective function (more in Appx. Sec. B).
Figure 4: An example of context-aware preference selection. We demonstrate our model’s ability to generate preference feedback, i.e., preferred choice and explanation marked in orange box, on given instruction input, prompt, two responses and context. Red color denotes an incorrect response, while blue color signifies a correctly grounded response with respect to the visual input. The rationale behind selecting ‘Output B’ as the preferred choice is indicated in green.
@@
We propose to explicitly integrate detailed video descriptions, referred to as context, into the preference selection workflow, thereby imparting additional contextual clarity to the VLMM, as illustrated in Figures 3-(2) and 4. Specifically, we start by segmenting the video into small clips, each containing up to 20 frames, and then employ the VLM-SFT to generate a detailed video description for each segment with an input prompt, ‘Describe this video in detail’. Subsequently, these individual captions are concatenated, which we call a narrative of the video. The narrative is then provided to a judge model, i.e., VLM-SFT, for better preference selection. The context not only improves the VLM-SFT’s ability to obtain a comprehensive view of the video content, but also enables it to identify the most suitable response for the video (see Sec. 4.3 for empirical results). Integrating the context with instruction inputs with a specific prompt (rules for generating preferences as illustrated in Appendix Fig. 9), marked in dotted boxes in Fig. 3-(2), allows us to collect context-aware preferences.
Training the reward model.
We design the reward model (RM) to assign higher scores to responses considered better and lower scores to those considered worse in quality. Starting from VLM-SFT with the final linear layer removed, we train a model to take an input prompt and response, i.e., marked in dotted boxes of Fig. 3-(3), and output a scalar reward value. Using the preference dataset produced from the VLM-SFT, we train the RM with a cross-entropy loss. Specifically, we use 13B VLM-SFT to train the reward model, as it gives slightly better performance than 7B VLM-SFT (see the quantitative comparison discussed in Sec. 4.3). Note that, after training RLAIF using the RM, the 7B VLM-RLAIF significantly surpass the 13B VLM-SFT, thus validating the effectiveness of our proposed framework in aligning video and text modalities (refer to Sec. 4.2 for more details).
3.2 Two-stage Curriculum SFT
During the SFT process, we initially train the LLM with a open-sourced video-text instruction-tune dataset Muhammad Maaz and Khan (2023). To improve the VLMM, we not only augment our training with additional video-text instruction-tune datasets but also propose a novel curriculum learning strategy.
@@ 3.2 Two-stage Curriculum SFT
To improve the video understanding ability, we first augment the video-text instruction-tune dataset Muhammad Maaz and Khan (2023) with existing human-annotated video question answering datasets Xiao et al. (2021); Li et al. (2020). In particular, we focus on obtaining instruction-tune dataset that encompass both visual and temporal intricacies for video comprehension. To obtaining such dataset, we use Video Localized Narratives Voigtlaender et al. (2023), a dataset that provides detailed descriptions of the appearance and complex actions of all objects in the video. We transform the narrative dataset into an object-centric instruction-tune dataset format using ChatGPT similar to prior work Muhammad Maaz and Khan (2023). More details about the generated instruction-tune dataset are in Appendix Sec. A.
Curriculum-based supervised fine-tuning.
Inspired by the human learning process, we argue the importance of the training sequence, the curriculum learning principles Chang et al. (2021) for learning VLMMs. We propose a simple two-stage curriculum learning strategy for SFT, aiming to optimize the instruction-following proficiency of VLMMs.
For the curriculum based structured learning, we divide the SFT dataset based on complexity into two segments: ‘easy’ and ‘hard’. The model is first trained on ‘easy’ tasks to learn basic concepts, then progressed to ‘hard’ tasks for advanced challenges. For the difficulty or easiness of the data, given that longer answers often require deeper comprehension of the context and enhanced proficiency in creating syntactically complex sentences Xu et al. (2020); Agrawal and Singh (2023); Ranaldi et al. (2023), we use answer length as our criterion for sample difficulty, i.e., the longer the answer sentence, the more difficult the task is.
@@
Correctness $\uparrow$ Detail $\uparrow$ Context $\uparrow$ Temporal $\uparrow$ Consistency $\uparrow$
VideoChat Li et al. (2023b) 7B 2.23 2.50 2.53 1.94 2.24
@@
Table 1: Quantitative comparison between different VLMMs on video-based generative performance benchmark. Our approach, VLM-RLAIF, shows a performance improvement over previous approaches, with the exception of GPT-4V which requires much more computational resource than ours, and demonstrates noticeable enhancements across five criteria when compared to the VLM-SFT. Here, $\Delta$ (RLAIF - SFT) indicates the improvement of RLAIF model over SFT model. $\dagger$} denotes reproduced results using the author’s implementation.